OpenStudy (anonymous):

list the first twenty counting numbers in base seven? (only digits 0 thru 6 are used in base 7)??????????????????????????????????????

OpenStudy (anonymous):

0, 1, 2, 3, 4, 5, 6 10, 11, 12, 13, 14, 15, 16 20, 21, 22, 23, 24, 25 is this right?

myininaya (myininaya):

1 2 3 4 5 6 7=7^1=10 8=7^1+7^0=11 9=7^1+2*7^0=12 10=7^1+3*7^0=13 11=7^1+4*7^0=14 12=7^1+5*7^0=15 13=7^1+6*7^0=16 14=2*7^1+0*7^0=20 15=2*7^1+1*7^0=21 16=2*7^1+2*7^0=22 17=2*7^1+3*7^0=23 18=2*7^1+4*7^0=24 19=2*7^1+5*7^0=25 20=2*7^1+6*7^0=26

myininaya (myininaya):

21=30 22=31 23=32 24=33 25=34 26=35 27=36 28=40 29=41 do you see a pattern?

OpenStudy (anonymous):

THE NUMBERS KEEP INCREASING THAT PATTERN?

OpenStudy (anonymous):

MYININAYA CAN U EXPLAIN FURTHER PLEASEE

myininaya (myininaya):

the 7^0 digit only stays between 0 and 6

OpenStudy (anonymous):

7/\1 MEANS 7 AND 1 CORRECT?

myininaya (myininaya):

7^0=1 7^1=7 7^2=49

myininaya (myininaya):

those are powers

myininaya (myininaya):

\[7^x\]=7^x

OpenStudy (anonymous):

YES OK I UNDERSTAND THAT NOW WE ARE TALKING OF BASE SEVEN CORRECT

OpenStudy (anonymous):

Just like in base 10 you have the ones column, the sevens column, the 49's column, etc. So if you take a number (like 56) It has 1 49 (with 17 left over) 2 7's (with 3 left over) and 3 ones. \(123_7 = 56_{10}\)

myininaya (myininaya):

say we are given a number c and we want to write it in base 7 first we would figure out the coefficients a_i of \[c=a_n7^n+a_{n-1}7^{n-1}+a_{n-2}7^{n-2}+\cdot \cdot \cdot + a_37^3+a_27^2+a_17^1+a_07^0\] then we can write c as \[c=a_na_{n-1}a_{n-2}....a_3a_2a_1a_0\] where he have written c in base 7

myininaya (myininaya):

where we have written c in base 7*

myininaya (myininaya):

\[c=a_na_{n-1}a_{n-2}....a_3a_2a_1a_0 \] this is not multiplication these are the digits of c

myininaya (myininaya):

in order of course

myininaya (myininaya):

maybe i should say \[c_{10}=(a_na_{n-1}a_{n-2}....a_3a_2a_1a_0)_{7}\]

OpenStudy (anonymous):

ust like in base 10 you have the ones column, the sevens column, the 49's column, etc. So if you take a number (like 56) It has 1 49 (with 17 left over) 2 7's (with 3 left over) and 3 ones. 1237=5610- I THINK I CAN UNDERSTAND THIS POINT

myininaya (myininaya):

\[56_{10}=1 \cdot 7^2+1 \cdot 7^1+0 \cdot 7^0=110_{7}\]

OpenStudy (anonymous):

ok, Myininaya u just lost me. your going to fast

OpenStudy (anonymous):

go back to my original question

OpenStudy (anonymous):

the answer is irrelevant. just tell me how u got it

OpenStudy (anonymous):

what are the counting numbers of a base?

myininaya (myininaya):

because 49+7=56

myininaya (myininaya):

say you want to write a number c in base 7 and you are given a number in base 10 you could do this to figure out what you highest power would be \[\log_7c=\frac{lnc}{\ln7}\] so for example we want to write 56 in base 7 so we do \[\log_7(56)=\frac{\ln(56)}{\ln7}=2.068621561\] take the floor of this number which is 2 and that is the highest power you should begin with so we have the 7^2 part \[56=1 \cdot 7^2+ a_1 \cdot 7^1+a_2 \cdot 7^0\] we need to figure out these other digits so we already have 49 and 56-49=7 so we should put a 1 in front of 7^1 to get the 7 and a 0 in front of 7^0 so we have \[56=1 \cdot 7^2+ 1 \cdot 7^1+0 \cdot 7^0=(110)_7\]

myininaya (myininaya):

\[\frac{56}{7}=8=7+1\] \[(\frac{56}{7}=7+1)7\] \[56=7(7)+1(7)=7^2+7\]

OpenStudy (anonymous):

1 2 3 4 5 6 7=7^1=10 8=7^1+7^0=11 9=7^1+2*7^0=12 10=7^1+3*7^0=13 11=7^1+4*7^0=14 12=7^1+5*7^0=15 13=7^1+6*7^0=16 14=2*7^1+0*7^0=20 15=2*7^1+1*7^0=21 16=2*7^1+2*7^0=22 17=2*7^1+3*7^0=23 18=2*7^1+4*7^0=24 19=2*7^1+5*7^0=25 20=2*7^1+6*7^0=26 explain this answer to me. how is 7 to the first power = to 10 how did u get the first six numbers

myininaya (myininaya):

10=7+3

OpenStudy (anonymous):

but 7 to the first power is not 7 + 3

myininaya (myininaya):

\[10=7^1+3 \cdot 7^0=(10)_7\]

myininaya (myininaya):

no 7^1 does not = 7+3 but 7^1+3*7^0=7+3

OpenStudy (anonymous):

this is what u have......7=7^1=10

myininaya (myininaya):

write u said to write it in base 7

myininaya (myininaya):

\[7=1 \cdot 7^1+0 \cdot 7^0=10_7\]

myininaya (myininaya):

right u said to write in base 7*

myininaya (myininaya):

\[1_{10}=1 \cdot 7^0=1_7\] the other first 5 work like this as well

myininaya (myininaya):

\[2_{10}=2 \cdot 7^0=2_{7}\]

myininaya (myininaya):

\[6_{10}=6 \cdot 7^0=6_7\]

OpenStudy (anonymous):

when i am asked to find the counting numbers in a base.......what is that?

myininaya (myininaya):

oh yes forgot to say one more thing you want all the a_i's to state between 0 and 7 not including 7

OpenStudy (anonymous):

HELP???????\[78000/1+0.08\left(\begin{matrix}17 \\ 12\end{matrix}\right)\]